RO VE and the region 2 2.-(50) Find Is, Ic of operation, 20012 ΙΜΩ #114 3V 61 200
Answer #1

imir JB 3V 2002 I Ic =200 ts IE IE Ict IB Ic=ß IB By KUL -3+ Ic (200)+ Ig 102) = 0 -3 + Iß (200)(200) + Ig (106) = 0 -3 + IBby KUL -3 + Ic (200) &VCE =D -3 -3 + 0.57X10°x200 + VCE = 0 -2.886 #VCE = Vce =9,886 V in between O and Vcc Here VCE is OL VC